OpenStudy (anonymous):

A jar contains 5 white balls, 4 red balls and 6 black balls. What is the probability that you will draw out a red or black ball?

OpenStudy (amistre64):

red or black dont matter so add up red and black balls; then divide by total

OpenStudy (anonymous):

Okay so the probability of drawing a red ball is 4/15 and the probability of drawing a black ball is 6/15. As it is an OR question, we use the OR rule: P(A) + P(B) - P(A+B) P(A) = 4/15, P(B) = 6/15 and P(A+B) doesn't exist because you can't draw a ball that is red AND black. So: 4/15 + 6/15 - 0 = 10/15 (which simplifies to 2/3)